### v1

@@ 3.3 Spectral Local Motion Refinement
Another problem in (17) is that the residuals between consecutive frames, while rich in motion information, may also encapsulate high-frequency local distortions, background noise, and other non-motion-related artifacts. Accordingly, we focus on particularly prioritizing low-to-moderate spatial frequency components. Specifically, following the amplitude and phase spectrum definition in (8), we define amplitude and phase matching loss as follows:
$\begin{split}&\ell_{\text{l}ocal}^{A}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))=\mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\||\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)|-|\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)|\|_{1}\Big{]},\\ &\ell_{\text{l}ocal}^{P}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))=\mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\|\angle\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)-\angle\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)\|_{1}\Big{]},\end{split}$ (20)
@@
$\omega(a,b)=\Big{[}\big{(}\frac{H}{2}\big{)}^{2}+\big{(}\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}-\Big{[}\big{(}a-\frac{H}{2}\big{)}^{2}+\big{(}b-\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}+1$ (21)
for $0<a<H,0<b<W$, and otherwise, set to zero. This introduces a weighting [40] that prioritizes low-frequency components for $\delta>0$.
@@ 3.4 Inference Pipeline
$\min_{\theta}\mathbb{E}_{t,n,{\boldsymbol{\epsilon}}_{t}^{n},{\boldsymbol{\epsilon}}_{t}^{n+1}}\Big{[}\ell_{\text{align}}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}+\lambda_{g}\ell_{\text{global}}\big{(}\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t)\big{)}+\lambda_{l}\ell_{\text{l}ocal}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}\Big{]},$ (22)
@@
This Spectral Motion Alignment (SMA) is universally adaptable across various motion distillation frameworks. While diverse diffusion-based motion distillation frameworks adopt their pixel-domain motion learning objectives, the proposed frequency-domain alignment seamlessly integrates with these arbitrary objectives. Moreover, different motion distillation frameworks target specific parameters $\theta$ for fine-tuning, varying from temporal attention layers [16] to dual-path LoRAs [49]. We empirically demonstrate global compatibility of the proposed spectral motion alignment with diverse neural architectures and parameterizations. Pseudo-code is provided in the appendix.
Figure 4: Comparison within VMC framework, using Show-1 cascaded model (top) and Zeroscope (bottom) non-cascaded model. Visit our project page for an in-depth view.
@@ 4.1 Experimental Setting
To assess the capability of Spectral Motion Alignment (SMA) to capture accurate motion within contemporary diffusion-based motion learning frameworks, we curated a dataset comprising 30 text-video pairs sourced from the publicly available DAVIS [24] and WebVid-10M [2] collections. This dataset is deliberately designed to cover a broad spectrum of motion types and subjects, with video lengths ranging between 8 and 16 frames. For this study, we leverage two foundational text-to-video diffusion models: Zeroscope [33], a non-cascaded VDM, and Show-1 [45], a cascaded VDM. More details are provided in appendix.
@@ 4.2 Baselines
4.2.1 MotionDirector.
MotionDirector [49] pioneered the concept of motion customization within video diffusion literature. To distinctively tailor the appearance and motion of a video, [49] developed a unique dual-path framework employing Low-Rank Adaptation (LoRA, [14]). Within this architecture, spatial LoRAs are optimized using the diffusion training loss on a single, randomly chosen frame, rather than on the full video sequence. Conversely, temporal LoRAs are trained by considering all frames of the video, using an anchor frame based appearance-debiased temporal loss.
4.2.2 VMC.
VMC [16] achieves state-of-the-art performance in motion customization through their novel epsilon residual matching objective, facilitating efficient motion distillation within cascaded video diffusion framework. In their work, they showcased the efficient of motion distillation within cascaded video diffusion models, by optimizing only keyframe generation model while freezing later modules. As VMC represents the most closely related work to ours, we conduct a two-fold validation within this baseline, including both cascaded and non-cascaded video diffusion scenarios.
4.2.3 Qualitative Comparison.
Fig. 3 offers a qualitative analysis comparing MotionDirector’s performance with and without the implementation of SMA. Likewise, Fig. 4 examines VMC’s performance with and without SMA implementation. Specifically, videos produced using a cascaded diffusion pipeline are displayed in Fig. 4-top, whereas those generated through a non-cascaded diffusion model are shown in the bottom. Videos generated without SMA correctly capture the appearance as dictated by the target text, but fail to accurately mirror the original video’s motion patterns. In contrast, SMA significantly enhances motion understanding, precisely distinguishing between dynamic and static elements in the scene. For instance, in the last example of Fig. 3, MotionDirector with SMA produces a video where only the eagle moves from right to left, whereas without SMA, the video inaccurately depicts the ground moving alongside the eagle.
4.2.4 Quantitative Comparison.
The results of our quantitative evaluation are presented in Table 1. We employ CLIP encoders [26] for automated evaluation. To evaluate text-video alignment [10], we measure the average cosine similarity between the target text prompt and the frames generated. Regarding frame consistency, we extract CLIP image features for each frame in the output video and subsequently calculate the average cosine similarity among all frame pairs in the video. For human evaluation, we conduct a user study with 42 participants to assess three key aspects, guided by the following questions: (1) Editing Accuracy: Is the output video accurately edited, reflecting the target text? (2) Temporal Consistency: Is the transition between frames smooth and consistent? (3) Motion Accuracy: Is the motion of the input video accurately preserved in the output video? Tab. 1 demonstrates that SMA enhances the performance of MotionDirector and VMC across all measured metrics.
@@
Figure 5: Comparison within Tune-A-Video and ControlVideo (Depth) baseline.
5 Experiments using T2I Video Diffusion Models
@@ 5.1 Experimental Setting
To assess the effectiveness of SMA within methodologies built on text-to-image diffusion model, we use the same text-video pairs previously assembled in Sec. 4.1. The resolution for all produced videos is standardized to 512x512. In this experiment, Stable Diffusion v1-5 [28] and ControlNet-Depth [46] are utilized.
@@ 5.2 Baselines
5.2.1 Tune-A-Video.
Tune-A-Video [38] was at the forefront of image-diffusion-based video editing at a time when video diffusion models were not yet publicly accessible. It begins with transforming a pretrained T2I diffusion model to psuedo T2V model by adding temporal attention layers and expanding spatial self-attention into spatio-temporal attention. Attention projection layers within this pseudo T2V model are then fine-tuned on a specific input video, utilizing the diffusion training objective (Eq. 2).
5.2.2 ControlVideo.
ControlVideo [48] is another one-shot-based video editing method, starting from pretrained T2I model. ControlVideo extends ControlNet [46] from image to video to incorporate structural cues obtained from the input video. Subsequently, akin to Tune-A-Video [38], attention projection layers of the backbone U-Net and ControlNet are customized on the input video.
@@
Fig. 5-(top) demonstrates the qualitative benefits of incorporating SMA into the Tune-A-Video method. Without SMA, Tune-A-Video often experiences flickering in foreground objects, a problem that SMA significantly reduces. Fig. 5-(bottom) illustrates the qualitative improvements brought by integrating SMA with the ControlVideo framework. In ControlVideo, depth control aids in maintaining the structural integrity from the input videos to the output videos. However, it’s observed that structural integrity alone doesn’t ensure motion accuracy, and SMA plays a crucial role in accurately capturing motion details.
@@
Quantitative results are detailed in Tab. 2, following the metrics introduced in Sec. 4.2.4. Across both the Tune-A-Video and ControlVideo frameworks, the integration of SMA results in enhanced performance across all five evaluated metrics, notably achieving a substantial advantage in motion accuracy.
Figure 6: Visualization of (a) spatial frequency spectrum and (b) motion vectors estimated from the pre-trained Show-1 [45] without fine-tuning. (c) Ablation study on spectral motion alignment based on VMC [16].
6 Discussion
We explore the impact of motion vector refinement by examining motion vectors ($\delta{\boldsymbol{v}}_{0}$, $\delta\hat{{\boldsymbol{v}}}_{0}(t)$) and their (amplitude, phase) spectrum in Figure 6 ($t=700$). Figure 6b indicates that the original motion vector $\delta{\boldsymbol{v}}_{0}$ filters out appearance (e.g. background, color, etc) to some extent. However, it still faces frame-wise distortions or inconsistencies. For instance, in a specific example of this skateboarding man, non-motion-related artifacts, e.g. stair and fence patterns, background texture, etc., persist as distortions. These are characterized as high-frequency noises in the amplitude spectrum, Fig. 6a. Conversely, $\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)$ shows a reduction in high-frequency distortions while preserving essential motion information. As shown in (14), $\delta{{\boldsymbol{v}}}_{t}^{n}$ can be acquired from $\delta\boldsymbol{v}_{0}^{n}$ through the diffusion kernel. In this forward process, high-frequency components are perturbed more rapidly [20]. Consequently, the core motion information is consistently preserved in $\delta\boldsymbol{v}_{t}^{n}$ from coarse ($t\uparrow$) to fine ($t\downarrow$) level. Motivated by these insights, we prioritize low spatial frequency components during motion distillation to avoid overfitting to non-motion-related high-frequency distortions.
This motion vector refinement improves the overall fidelity as shown in Figure 6c. The application of VMC for motion transfer, without refinement, often fails to maintain motion patterns, introducing repetitive fence-like background distortions. Conversely, motion refinement within the frequency domain significantly improves fidelity and background detail. Moreover, global motion alignment further improves motion transfer. Specifically, while local motion alignment is generally effective, it occasionally generates "reversed" motions, i.e. an astronaut skateboarding in an upward direction. Note that this example is particularly challenging as it shows very subtle positional changes (video in project page). This highlights the limitations of conventional frameworks in identifying accurate motion from a single motion vector frame $\delta{\boldsymbol{v}}_{0}^{n}$, Figure 6b. In contrast, the proposed global motion alignment effectively mitigates these challenges, ensuring accurate learning of motion patterns. Fig. 7 further demonstrates the effectiveness of global motion alignment. Thus, our approach harmonizes both local and global motion alignment and effectively distills motion patterns while robustly reducing high-frequency distortions.
Figure 7: Ablation study on $\ell_{\text{global}}$. Global motion alignment facilitates motion transfer.
@@ 7 Conclusion
In this paper, we propose Spectral Motion Alignment (SMA), a novel frequency-based distillation framework for motion transfer. We explore the limitations of conventional motion estimation methods: (a) lack of global motion understanding, (b) vulnerability to spatial artifacts. By utilizing Fourier and wavelet transforms, SMA captures comprehensive motion dynamics, highlighting the significance of global motion context. We proved SMA’s effectiveness and compatibility through extensive validation across different motion transfer scenarios.
Reproducibility & Ethics Statement. For reproducibility, the code and datasets will be released to the public through the project page. Video generative models may pose potential risks for misuse in creating harmful content, which necessitates political regulation.
Appendix 0.A Pseudo Training Algorithm
In our work, we adopt the notation and expressions mostly from [16] for the preliminaries section and pseudo-code, due to its relevance to our focus on denoised motion vector estimates. We interchangeably use $\hat{{\boldsymbol{v}}}_{0}^{1:N}(t)$ and $\hat{{\boldsymbol{v}}}_{0}(t)$ in the main paper and Algorithm 1. While Algorithm 1 generalizes parameter $\theta$, each video diffusion model incorporates specific parameters such as $\theta_{\text{TA}}$ [16] and $\theta_{\text{LoRA}}$ [49].

### v2

@@ 3.3 Spectral Local Motion Refinement
Another problem in (14) is that the estimated motion representations may encapsulate high-frequency local distortions, background noise, and other non-motion-related artifacts. By aligning the denoised estimates with these artifacts, the fine-tuned VDM may erroneously reproduce similar high-frequency artifacts as in Figure 6.
Accordingly, we focus on prioritizing low-to-moderate spatial frequency components particularly. Specifically, following the amplitude and phase spectrum definition in (6), we define amplitude and phase matching loss, $\ell_{\text{l}ocal}^{A}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))$ and $\ell_{\text{l}ocal}^{P}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))$, as follows:
$\begin{split}&(A)\ \mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\||\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)|-|\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)|\|_{1}\Big{]},\\ &(P)\ \mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\|\angle\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)-\angle\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)\|_{1}\Big{]},\end{split}$ (17)
@@
$\begin{split}\omega(a,b)=\Big{[}\big{(}\frac{H}{2}\big{)}^{2}+\big{(}\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}-\Big{[}\big{(}a-\frac{H}{2}\big{)}^{2}+\big{(}b-\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}+1\end{split}$
for $0<a<H,0<b<W$, and otherwise, set to zero. This introduces a weighting (Yang et al. 2022) that prioritizes low-frequency components for $\delta>0$.
@@ 3.4 Inference Pipeline
$\begin{split}\min_{\theta}&\mathbb{E}_{t,n,{\boldsymbol{\epsilon}}_{t}^{n},{\boldsymbol{\epsilon}}_{t}^{n+1}}\Big{[}\ell_{\text{align}}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}+\\ &\lambda_{g}\ell_{\text{global}}\big{(}\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t)\big{)}+\lambda_{l}\ell_{\text{l}ocal}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}\Big{]},\end{split}$ (18)
@@
This Spectral Motion Alignment (SMA) is universally adaptable across various motion distillation frameworks. While diverse diffusion-based motion distillation frameworks adopt their pixel-domain motion learning objectives, the proposed frequency-domain alignment seamlessly integrates with these arbitrary objectives. Moreover, different motion distillation frameworks target specific parameters $\theta$ for fine-tuning, varying from temporal attention layers (Jeong, Park, and Ye 2023) to dual-path LoRAs (Zhao et al. 2023b). We empirically demonstrate the global compatibility of the proposed spectral motion alignment with diverse neural architectures and parameterizations. Pseudo-code is provided in the appendix.
3.5 Extending SMA to Feature Space
Beyond pixel-space motion representations, SMA can be further extended to semantic diffusion features (DIFT, (Tang et al. 2023)). Specifically, Diffusion-Motion-Transfer (DMT, (Yatim et al. 2023)) constructs motion vectors based on pairwise differences in space-time diffusion features, which are then utilized for latent optimization-based video motion transfer. Given input and target video latents, ${\boldsymbol{v}}_{t}$ and $\tilde{{\boldsymbol{v}}}_{t}$, the model extracts space-time features $f({\boldsymbol{v}}_{t})$ and $f(\tilde{{\boldsymbol{v}}}_{t})$. Then, the feature residuals are defined as $\delta f({\boldsymbol{v}}_{t})^{n}$ and $\delta f(\tilde{{\boldsymbol{v}}}_{t})^{n}$, the $n$-th consecutive difference between hidden feature frames. This leads to the spectral alignment objective in feature space as follows:
$\begin{split}\mathbb{E}\Big{[}&\ell_{\text{DMT}}\big{(}f({\boldsymbol{v}}_{t}),f(\tilde{{\boldsymbol{v}}}_{t})\big{)}+\\ &\lambda_{g}\ell_{\text{global}}\big{(}\delta f({\boldsymbol{v}}_{t}),\delta f(\tilde{{\boldsymbol{v}}}_{t})\big{)}+\lambda_{l}\ell_{\text{l}ocal}\big{(}\delta f({\boldsymbol{v}}_{t})^{n},\delta f(\tilde{{\boldsymbol{v}}}_{t})^{n}\big{)}\Big{]},\end{split}$ (19)
where $\ell_{\text{DMT}}$ refers to the original space-time feature loss in (Yatim et al. 2023). Note that DMT does not finetune the models, leveraging (19) for a latent optimization in sampling process. We demonstrate the effectiveness of spectral alignment in the diffusion feature space by comparing it against the original DMT framework in Fig.4-bottom.
Figure 4: Comparison within VMC framework using Show-1 video model (top) and DMT framework using Zeroscope video model (bottom). Each demonstrate the compatibility of SMA in pixel-space and feature-space, respectively.
@@ 4.1 Experimental Setting
To assess the capability of Spectral Motion Alignment (SMA) to capture accurate motion within contemporary diffusion-based motion learning frameworks, we curated a dataset comprising 30 text-video pairs sourced from the publicly available DAVIS (Pont-Tuset et al. 2017) and WebVid-10M (Bain et al. 2021) collections. This dataset is deliberately designed to cover a broad spectrum of motion types and subjects, with video lengths ranging between 8 and 16 frames. For this study, we leverage two foundational text-to-video diffusion models: Zeroscope (Sterling 2023), a non-cascaded VDM, and Show-1 (Zhang et al. 2023a), a cascaded VDM. More details are provided in appendix.
@@ 4.2 Baselines
MotionDirector (Zhao et al. 2023b) tailor the appearance and motion of a video by developing a unique dual-path (spatial, temporal) framework with Low-Rank Adaptation (LoRA, (Hu et al. 2021)). VMC (Jeong, Park, and Ye 2023) achieves state-of-the-art performance in motion customization through their novel epsilon residual matching objective, facilitating efficient motion distillation within a cascaded video diffusion. DMT (Yatim et al. 2023) proposes a new space-time feature loss, guiding the sampling process towards preserving the motion patterns while complying with the target object. Please see Sec 3.5 for more details.
4.3 Qualitative Comparison.
Fig. 3 and 4 offer qualitative comparisons with and without SMA. The top of Figure 4 shows videos from a cascaded diffusion pipeline, while the bottom displays those from a non-cascaded model. Without SMA, videos may capture appearance to some extent but fail to replicate motion patterns accurately. In contrast, SMA significantly improves motion transfer, distinguishing dynamic from static objects. For instance, in the last example of Fig. 3, SMA produces a video where only the eagle moves from right to left, whereas without SMA, the video inaccurately depicts the ground moving alongside the eagle.
4.4 Quantitative Comparison.
The results of our quantitative evaluation are presented in Table 1. To evaluate text-video alignment (Hessel et al. 2021), we measure the average cosine similarity between the target text prompt and the frames generated. Regarding frame consistency, we extract CLIP image features for each frame in the output video and subsequently calculate the average cosine similarity among all frame pairs in the video. For human evaluation, we conduct a user study with 42 participants to assess three key aspects, guided by the following questions: (1) Editing Accuracy: Is the output video accurately edited, reflecting the target text? (2) Temporal Consistency: Is the transition between frames smooth and consistent? (3) Motion Accuracy: Is the motion of the input video accurately preserved in the output video? Tab. 1 demonstrates that SMA enhances the performance of MotionDirector and VMC across all measured metrics.
@@
Figure 5: Comparison within Tune-A-Video (Top) and ControlVideo-Depth (Bottom) baseline.
Figure 6: Visualization of (a) spatial frequency spectrum and (b) motion vectors estimated from the pre-trained Show-1 (Zhang et al. 2023a) without fine-tuning. (c) Ablation study on spectral motion alignment based on VMC (Jeong, Park, and Ye 2023).
5 T2I Video Diffusion Models
@@ 5.1 Experimental Setting
We further evaluate the efficacy of SMA with methods based on text-to-image diffusion model. Same text-video pairs are used as in Sec. 4.1. The resolution for all produced videos is standardized to 512x512. In this experiment, Stable Diffusion v1-5 (Rombach et al. 2022) and ControlNet-Depth (Zhang, Rao, and Agrawala 2023) are utilized.
@@ 5.2 Baselines
Tune-A-Video (Wu et al. 2023) transforms a pretrained T2I diffusion model to psuedo T2V model by adding temporal attention layers and expanding spatial self-attention into spatio-temporal attention. ControlVideo (Zhao et al. 2023a) is another one-shot-based video editing method stems from pretrained T2I model. ControlVideo extends ControlNet (Zhang, Rao, and Agrawala 2023) from image to video to incorporate structural cues obtained from the input video.
@@
Fig. 5-(top) demonstrates the efficacy of SMA with Tune-A-Video method, where SMA alleviates the flickering artifacts in foreground objects. Fig. 5-(bottom) further illustrates the improvements with the ControlVideo framework. While depth control encourages ControlVideo to maintain the structural integrity, Fig. 5 shows that it is not sufficient for motion accuracy, where SMA plays a crucial role in accurate capture of motion details.
@@
Quantitative results are detailed in Tab. 2, following the metrics introduced in Sec. 4.4. Across both the Tune-A-Video and ControlVideo frameworks, the integration of SMA improves performance across all five evaluated metrics, notably achieving a substantial advantage in motion accuracy.
6 Analysis
We explore the impact of SMA by examining motion vectors ($\delta{\boldsymbol{v}}_{0}$, $\delta\hat{{\boldsymbol{v}}}_{0}(t)$) and their (amplitude, phase) spectrum in Figure 6 ($t=700$). Figure 6b indicates that the pixel-space motion vector $\delta{\boldsymbol{v}}_{0}$ faces frame-wise distortions or inconsistencies. For instance, the motion-independent artifacts, e.g. stair and fence patterns, background texture, etc., persist as distortions in the obtained motion vectors. These are characterized as high-frequency noises in the amplitude spectrum, Fig. 6a. Motivated by these observations, we prioritize low spatial frequency components during motion distillation to avoid overfitting to motion-independent high-frequency distortions. This representations refinement improves the overall fidelity and removes background distortions (Figure 6c).
Moreover, global motion alignment further improves motion transfer. Specifically, without considering the global motion dynamics, existing frameworks occasionally generate “reversed" motions, i.e. an astronaut skateboarding in an upward direction. This highlights the limitations of conventional frameworks in understanding accurate motion from a single frame of motion vector $\delta{\boldsymbol{v}}_{0}^{n}$. In contrast, the proposed global motion alignment effectively mitigates these challenges, ensuring accurate learning of motion patterns. Table 3 demonstrates the importance of both global/local terms (More qualitative ablation results in appendix).
Baseline Baseline $+\mathcal{L}_{\text{local}}$ Baseline $+\mathcal{L}_{\text{global}}$ Baseline $+\mathcal{L}_{\text{local}}+\mathcal{L}_{\text{global}}$
Text-Align 0.8163 0.8382 0.8344 0.8427 (+ 0.026)
Frame-Con 0.9497 0.9613 0.9585 0.9659 (+ 0.016)
Table 3: Quantitative ablation of $\mathcal{L}_{\text{local}}$ and $\mathcal{L}_{\text{global}}$.
@@ 7 Conclusion
We propose Spectral Motion Alignment (SMA), a novel motion distillation framework in spectral domain. We explore the limitations of conventional motion estimation methods: (a) lack of global motion understanding, (b) vulnerability to spatial artifacts. Then, we mitigate these problems by harmonizing both local and global motion alignment and effectively distills motion patterns.
Acknowledgments
This work was supported by Institute of Information &\&& communications Technology Planning &\&& Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program(KAIST), No. RS-2023-00233251, System3 reinforcement learning with high-level brain functions), National Research foundation of Korea(NRF) (**RS-2023-00262527**, RS-2024-00336454, RS-2024-00341805).
Appendix A Pseudo Training Algorithm
In our work, we adopt the notation and expressions mostly from (Jeong, Park, and Ye 2023) for the preliminaries section and pseudo-code, due to its relevance to our focus on denoised motion vector estimates. We interchangeably use $\hat{{\boldsymbol{v}}}_{0}^{1:N}(t)$ and $\hat{{\boldsymbol{v}}}_{0}(t)$ in the main paper and Algorithm 1. While Algorithm 1 generalizes parameter $\theta$, each video diffusion model incorporates specific parameters such as $\theta_{\text{TA}}$ (Jeong, Park, and Ye 2023) and $\theta_{\text{LoRA}}$ (Zhao et al. 2023b).